OpenStudy (anonymous):

Prove: 2+2=5 Do your best to explain this even though we all know 2+2=4

OpenStudy (tfht):

its simply impossible

OpenStudy (anonymous):

Nope not true

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

2+2 = 2+2+0 correct? right. Now according to Robins Lemma proof Let x be a non-zero number, and set y=x. Thus: x = y x^2 = xy x^2-y^2 = xy-y^2 (x+y)(x-y) = y(x-y) x+y = y 2y = y 2 = 1 1 = 0 Now since 2+2 = 2+2+0 and we apply the lemma proof so that 0 = 1, 2+2+0 = 2+2+1 =5 THEREFORE 2+2 = 5

OpenStudy (anonymous):

can not divide by 0. if x = y, then x-y = 0

OpenStudy (anonymous):

\[2/0=\infty\]